# cuPDLP+: A Further Enhanced GPU-Based First-Order Solver for Linear Programming

### v1

@@ Abstract
We introduce cuPDLP+, a further enhanced GPU-based first-order solver for linear programming. Building on the predecessor cuPDLP, cuPDLP+ incorporates recent algorithmic advances, including the restarted Halpern PDHG method with reflection, a novel restart criterion, and a PID-controlled primal weight update. These innovations are carefully tailored for GPU architectures and deliver substantial empirical gains. On a comprehensive benchmark of MIPLIB LP relaxations, cuPDLP+ achieves 2x – 4x speedup over cuPDLP, with particularly strong improvements in high-accuracy and presolve-enabled settings.
@@ 1 Introduction
Recently, first-order methods (FOMs) [16] have emerged as a compelling alternative for solving large-scale LPs, offering a new paradigm that complements the traditional dominance of simplex and interior-point methods. Unlike simplex and interior-point methods, which rely on matrix factorizations and sequential computations, FOMs operate using simple iterative updates, primarily matrix-vector multiplications, making them particularly attractive for large-scale LPs. A notable breakthrough in this space is PDLP [1, 2], a solver based on the primal-dual hybrid gradient method (PDHG) [5] and tailored specifically for LP. By incorporating various practical enhancements, PDLP achieves significantly improved numerical stability and convergence in practice, positioning it as a practical and scalable solver for large-scale LPs within the FOM framework.
@@
Particularly, cuPDLP [12, 14] is a GPU-accelerated solver that extends PDLP by adapting its first-order framework to modern GPU architectures, with several algorithmic changes to make it more suitable to GPU architecture. The initial implementation, developed in Julia and referred to as cuPDLP [12], offloads key linear algebra operations, such as sparse matrix-vector multiplications, to the GPU and incorporates GPU-friendly heuristics, achieving notable speedups on medium-to-large-scale LP instances. Serving as both a practical solver and a research platform, cuPDLP illustrates how careful alignment between algorithm design and hardware capabilities can yield substantial performance improvements. A subsequent C implementation, cuPDLP-C [14], was developed to facilitate integration with production-scale computing environments. These advances in GPU-based LP solvers have attracted strong interest from both optimization software companies and technology firms, and have influenced the design of several commercial solvers, including Gurobi, COPT, FICO Xpress and NVIDIA’s cuOpt.
@@
These recent GPU-based solvers demonstrate the growing interest and practical potential of first-order methods when carefully adapted to modern hardware. Building on these foundations, we present cuPDLP+, a further enhanced GPU-accelerated linear programming solver. Compared to its predecessor cuPDLP, cuPDLP+ introduces several fundamental algorithmic changes that are primarily motivated by recent theoretical developments.
  • While cuPDLP is based on the restarted averaged PDHG (raPDHG) method [3], cuPDLP+ adopts the restarted Halpern PDHG (rHPDHG), a refinement inspired by recent theoretical insights[13]. This change allows the algorithm to take more aggressive steps, resulting in stronger empirical performance. The observed numerical improvements of cuPDLP+ over cuPDLP highlight the practical advantages of the rHPDHG framework.
  • In addition, cuPDLP+ explores a distinct set of enhancements and heuristics that further accelerate performance. These include a constant step-size rule, a restart condition aligned with the theoretical guarantees of rHPDHG [13], and a novel PID controller for updating the primal weight.
These algorithmic improvement translate into strong empirical performance of cuPDLP+. In various scenarios, cuPDLP+ achieves a 2× to 4× speedup. For example, in the high-accuracy setting with presolve, cuPDLP+ achieves a 2.9× speedup on 383 MIPLIB instances and a 4.63× speedup on hard instances, compared to cuPDLP.
@@ 1.1 Paper organization
Section 2 briefly introduces the form of LP to solve and the base algorithm vanilla PDHG. In Section 3, several algorithmic enhancement on top of cuPDLP are proposed to accelerate convergence. The numerical comparisons between cuPDLP+ and cuPDLP are presented in Section 4.
@@ 1.2 Notations
For a symmetric positive semidefinite matrix $M\in\mathbb{R}^{n\times n}$, the $M$-norm of a vector $x\in\mathbb{R}^{n}$ is defined as $\|x\|_{M}:=\sqrt{x^{\top}Mx}$. The orthogonal projector $\operatorname{proj}_{C}(x)$ denotes the Euclidean projection of a point $x\in\mathbb{R}^{n}$ onto a closed convex set $C\subseteq\mathbb{R}^{n}$, i.e., $\operatorname{proj}_{C}(x):=\arg\min_{y\in C}\|x-y\|_{2}$. For a set $C\subseteq\mathbb{R}^{n}$, we define $-C:=\{-x:x\in C\}$ as the reflection of $C$ about the origin. The $(i,j)$-th entry for a matrix $M\in\mathbb{R}^{m\times n}$ is denoted by $(M)_{ij}$. Given two vectors ({}$\ell\in(\mathbb{R}\cup\{-\infty\})^{n}$ and $u\in(\mathbb{R}\cup\{\infty\})^{n}$ satisfying $\ell\leq u$, define the mapping $p:\mathbb{R}^{n}\times(\mathbb{R}\cup\{-\infty\})^{n}\times(\mathbb{R}\cup\{\infty\})^{n}\to\mathbb{R}\cup\{\infty\}$ given by $p(y;\ell,u):=u^{\top}y^{+}-\ell^{\top}y^{-}$.
@@ 2 Preliminaries
In this section, we introduce the LP form that cuPDLP+ solves, followed by discussion on vanilla PDHG for solving LPs.
@@ 2.1 Linear programming
cuPDLP+ solves LP with the following primal-dual form:
$\min_{x\in\mathcal{X}}~{}~{}&~{}c^{\top}x&\\ \text{subject to:}~{}~{}&~{}Ax\in\mathcal{S}\ ,\end{aligned}\qquad\qquad\qquad\begin{aligned} \max_{y\in\mathcal{Y},r\in\mathcal{R}}\quad&-p(-y;\ell_{c},u_{c})-p(-r;\ell_{v},u_{v})\\ \text{subject to:}\quad&c-A^{\top}y=r\ ,\end{aligned}$ (1)
@@
$&\mathcal{Y}_{i}:=\begin{cases}\{0\}&(\ell_{c})_{i}=-\infty,~{}(u_{c})_{i}=\infty,\\ {\mathbb{R}}^{-}&(\ell_{c})_{i}=-\infty,~{}(u_{c})_{i}\in{\mathbb{R}},\\ {\mathbb{R}}^{+}&(\ell_{c})_{i}\in{\mathbb{R}},~{}(u_{c})_{i}=\infty,\\ {\mathbb{R}}&\text{otherwise};\end{cases}\quad\text{ and }\qquad\end{aligned}\begin{aligned} &\mathcal{R}_{i}:=\begin{cases}\{0\}&(\ell_{v})_{i}=-\infty,~{}(u_{v})_{i}=\infty,\\ {\mathbb{R}}^{-}&(\ell_{v})_{i}=-\infty,~{}(u_{v})_{i}\in{\mathbb{R}},\\ {\mathbb{R}}^{+}&(\ell_{v})_{i}\in{\mathbb{R}},~{}(u_{v})_{i}=\infty,\\ {\mathbb{R}}&\text{otherwise};\end{cases}\ .\end{aligned}$
@@ 2.2 PDHG
PDHG serves as the base routine of cuPDLP+ to solve primal-dual problem (2). Specifically, the update rule of PDHG on (2) is given as:
@@
where $\sigma$ and $\tau$ are dual and primal stepsizes respectively. cuPDLP+ further reparameterizes as $\tau=\frac{\eta}{\omega}$ and $\sigma=\eta\omega$, where $\eta$ is called the stepsize and $\omega$ is the primal weight. And the canonical norm of PDHG is defined by $\|\cdot\|_{P}$, where $P:=P_{\eta,\omega}=\begin{bmatrix}\frac{\omega}{\eta}I&A^{\top}\\ A&\frac{1}{\eta\omega}I\end{bmatrix}$. For notational convenience, we define the primal-dual iterate at iteration $k$ as $z^{k}=\begin{bmatrix}x^{k}\ y^{k}\end{bmatrix}$, and write $z^{k+1}=\operatorname{PDHG}(z^{k})$ to denote a single PDHG update step applied to $z^{k}$ as defined in (3).
@@ 3 Algorithmic enhancements on top of cuPDLP
To improve the efficiency and convergence behavior of cuPDLP+, we introduce a series of algorithmic enhancements on top of cuPDLP , inspired by recent advances in first-order methods [13]. These enhancements, implemented in our further enhanced solver cuPDLP+ , build on the PDHG framework and include techniques such as the Halpern iteration scheme, reflected updates, adaptive restarting, stepsize selection and dynamic primal-dual weight. Each component is carefully designed to accelerate convergence and enhance stability for solving linear programming problems. We detail these contributions below.
Halpern scheme. The Halpern scheme and its reflected variant are recent enhancements to the PDHG algorithm, aimed at improving convergence properties and accelerating solver performance. Originally developed to accelerate general operator splitting methods, the Halpern scheme has been successfully adapted to linear programming such as HPR-LP [6] and MPAX [11].
Halpern PDHG interpolates between the current PDHG iterate and the initial point, using a weighted average. Specifically, the update rule at iteration $k$ is:
@@
Reflection. The reflection technique builds upon the Halpern scheme by applying a reflected version of the PDHG operator, $(1+\gamma)\operatorname{PDHG}-\gamma\operatorname{id}$ with reflection $\gamma\in[0,1]$, instead of the vanilla PDHG update. This leads to the Reflected Halpern PDHG update:
@@
The reflection mechanism effectively takes a longer step compared to the vanilla Halpern update and has been shown to improve convergence both in theory and in practice [13].
Adaptive restart. Restarting is a key enhancement on first-order methods for attaining high-accuracy solutions. A restart is triggered when certain progress criteria are met, and the algorithm restarts from a new initial solution. In the Halpern scheme, the anchor is periodically reset to the current solution. This helps the algorithm stay focused on the neighborhood of the optimal solution, especially as the initial anchor becomes increasingly outdated during iterations. Such restart strategies have been shown to accelerate convergence rates for Halpern PDHG [13].
cuPDLP+ adopts an adaptive restart strategy that evaluates potential restart conditions at each iteration. The core idea is to monitor a fixed-point error metric $r(z)=\|z-\operatorname{PDHG}(z)\|_{P}$ at solution $z$, and trigger a restart when specific decay patterns are observed. This fixed-point error metric is motivated by the recent theoretical insight [13], and it is different from the normalized duality gap used in CPU-based PDLP[1, 2] and the KKT error used in cuPDLP[12]. In addition, three restart conditions are used:
@@
Stepsize. In contrast to an adaptive stepsize heuristic used in cuPDLP, cuPDLP+ switches to constant stepsize $\eta=\frac{0.99}{\|A\|_{2}}$, where operator norm of constraint matrix $\|A\|_{2}$ is approximated by power iteration.
Primal weight update. The primal weight $\omega$ is designed to balance the progress in the primal and dual spaces by adjusting its value. Specifically, the intuition is to determine the primal weight $\omega^{n}$ such that the weighted distances to optimality in the primal and dual domains are in the same scale, i.e., $\|(x^{n,t}-x^{*}),0\|_{w^{n}}\approx\|(0,y^{n,t}-y^{*})\|_{w^{n}}$. However, the optimal solution $(x^{,}y^{)}$ is not known during the iterations. To address this, cuPDLP estimates the primal and dual distances to optimality based on the observed movements in the previous epoch. The distance estimates are smoothed using exponential averaging and subsequently used to update the primal weight at each restart.
In cuPDLP+, we further enhance this strategy by modeling the update as a control problem and introducing a PID controller to dynamically regulate the primal weight. We define the error as the gap between the primal and dual distances on a logarithmic scale:
$e^{n}=\log\left(\frac{\sqrt{w^{n}}\left\|x^{n,t}-x^{*}\right\|_{2}}{\frac{1}{\sqrt{w^{n}}}\left\|y^{n,t}-y^{*}\right\|_{2}}\right)$

### v4

# cuPDLPx: A Further Enhanced GPU-Based First-Order Solver for Linear Programming

@@ Abstract
We introduce cuPDLPx , a further enhanced GPU-based first-order solver for linear programming. Building on the recently developed restarted Halpern PDHG for LP [16], cuPDLPx incorporates a number of new techniques, including a new restart criterion and a PID-controlled primal weight update. These improvements are carefully tailored for GPU architectures and deliver substantial computational gains. Across benchmark datasets, cuPDLPx achieves 2.5×–5× speedups on MIPLIB LP relaxations and 3×–6.8× on Mittelmann’s benchmark set, with particularly strong improvements in high-accuracy and presolve-enabled settings. The solver is publicly available at https://github.com/MIT-Lu-Lab/cuPDLPx.
@@ 1 Introduction
Recently, first-order methods (FOMs) [20] have emerged as a compelling alternative for solving large-scale LPs, offering a new paradigm that complements the traditional dominance of simplex and interior-point methods. Unlike simplex and interior-point methods, which rely on matrix factorizations and sequential computations, FOMs operate using simple iterative updates, primarily matrix-vector multiplications, making them particularly attractive for large-scale LPs. A notable breakthrough in this space is PDLP [4, 5], a solver based on the restarted average primal-dual hybrid gradient method (PDHG) [6, 8] and tailored specifically for LP. By incorporating various practical enhancements, PDLP achieves significantly improved numerical stability and convergence in practice, positioning it as a practical and scalable solver for large-scale LPs within the FOM framework.
@@
Particularly, cuPDLP [15, 17] is a GPU-accelerated solver that extends PDLP by adapting its first-order framework to modern GPU architectures, with several algorithmic changes to make it more suitable for GPU architectures. The initial implementation, developed in Julia and referred to as cuPDLP [15], offloads key linear algebra operations, such as sparse matrix-vector multiplications, to the GPU and incorporates GPU-friendly heuristics, achieving notable speedups on medium-to-large-scale LP instances. Serving as both a practical solver and a research platform, cuPDLP illustrates how careful alignment between algorithm design and hardware capabilities can yield substantial performance improvements. A subsequent C implementation, cuPDLP-C [17], was developed to facilitate integration with production-scale computing environments. These advances in GPU-based LP solvers have attracted strong interest from both optimization software companies and technology firms, and have influenced the design of several commercial solvers, including Gurobi, COPT, FICO Xpress and NVIDIA’s cuOpt.
@@
These recent GPU-based solvers demonstrate the growing interest and practical potential of first-order methods when carefully adapted to modern hardware. Building on these foundations, we present cuPDLPx , a further enhanced GPU-accelerated linear programming solver. Compared to its predecessor cuPDLP, cuPDLPx introduces several fundamental algorithmic changes that are primarily motivated by recent theoretical developments.
  • While cuPDLP is based on the restarted averaged PDHG (raPDHG) method [6], cuPDLPx adopts the restarted Halpern PDHG (rHPDHG), a refinement inspired by recent theoretical insights[16]. This change allows the algorithm to take more aggressive steps, resulting in stronger empirical performance. The observed numerical improvements of cuPDLPx over cuPDLP highlight the practical advantages of the rHPDHG framework.
  • In addition, cuPDLPx explores a distinct set of enhancements and heuristics that further accelerate performance. These include a constant step-size rule, a restart condition aligned with the theoretical guarantees of rHPDHG [16], and a novel PID controller for updating the primal weight.
In addition to these algorithmic advances, cuPDLPx is developed in C, providing a lightweight and highly efficient implementation that complements the earlier Julia prototype. These improvements together translate into strong empirical performance of cuPDLPx . On the MIPLIB relaxations, cuPDLPx achieves a 2.5×–5× speedup over cuPDLP, while on Mittelmann’s benchmark set, the gains are even larger, reaching 3×–6.8× speedups across different tolerances and presolve settings.
The solver is publicly available at https://github.com/MIT-Lu-Lab/cuPDLPx.
@@ 1.1 Paper organization
Section 2 briefly introduces the form of LP to solve and the base algorithm, vanilla PDHG. In Section 3, several algorithmic enhancements on top of cuPDLP are proposed to accelerate convergence. The numerical comparisons between cuPDLPx and cuPDLP are presented in Section 4.
@@ 1.2 Notations
For a symmetric positive semidefinite matrix $M\in\mathbb{R}^{n\times n}$, the $M$-norm of a vector $x\in\mathbb{R}^{n}$ is defined as $\|x\|_{M}:=\sqrt{x^{\top}Mx}$. The orthogonal projector $\operatorname{proj}_{C}(x)$ denotes the Euclidean projection of a point $x\in\mathbb{R}^{n}$ onto a closed convex set $C\subseteq\mathbb{R}^{n}$, i.e., $\operatorname{proj}_{C}(x):=\arg\min_{y\in C}\|x-y\|_{2}$. For a set $C\subseteq\mathbb{R}^{n}$, we define $C:=\{-x:x\in C\}$ as the reflection of $C$ about the origin. The $(i,j)$-th entry for a matrix $M\in\mathbb{R}^{m\times n}$ is denoted by $(M)_{ij}$. Given two vectors $\ell\in(\mathbb{R}\cup\{-\infty\})^{n}$ and $u\in(\mathbb{R}\cup\{\infty\})^{n}$ satisfying $\ell\leq u$, define the mapping $p:\mathbb{R}^{n}\times(\mathbb{R}\cup\{-\infty\})^{n}\times(\mathbb{R}\cup\{\infty\})^{n}\to\mathbb{R}\cup\{\infty\}$ given by $p(y;\ell,u):=u^{\top}y^{+}-\ell^{\top}y^{-}$.
@@ 2 Preliminaries
In this section, we introduce the LP form that cuPDLPx solves, followed by discussions on vanilla PDHG for solving LPs.
@@ 2.1 Linear programming
cuPDLPx solves LP with the following primal-dual form:
$\begin{aligned} \min_{x\in\mathcal{X}}~~&~c^{\top}x&\\ \text{subject to:}~~&~Ax\in\mathcal{S}\ ,\end{aligned}\qquad\qquad\qquad\begin{aligned} \max_{y\in\mathcal{Y},r\in\mathcal{R}}\quad&-p(-y;\ell_{c},u_{c})-p(-r;\ell_{v},u_{v})\\ \text{subject to:}\quad&c-A^{\top}y=r\ ,\end{aligned}$ (1)
@@
$\begin{aligned} &\mathcal{Y}_{i}:=\begin{cases}\{0\}&(\ell_{c})_{i}=-\infty,~(u_{c})_{i}=\infty,\\ {\mathbb{R}}^{-}&(\ell_{c})_{i}=-\infty,~(u_{c})_{i}\in{\mathbb{R}},\\ {\mathbb{R}}^{+}&(\ell_{c})_{i}\in{\mathbb{R}},~(u_{c})_{i}=\infty,\\ {\mathbb{R}}&\text{otherwise};\end{cases}\quad\text{ and }\qquad\end{aligned}\begin{aligned} &\mathcal{R}_{i}:=\begin{cases}\{0\}&(\ell_{v})_{i}=-\infty,~(u_{v})_{i}=\infty,\\ {\mathbb{R}}^{-}&(\ell_{v})_{i}=-\infty,~(u_{v})_{i}\in{\mathbb{R}},\\ {\mathbb{R}}^{+}&(\ell_{v})_{i}\in{\mathbb{R}},~(u_{v})_{i}=\infty,\\ {\mathbb{R}}&\text{otherwise};\end{cases}\ .\end{aligned}$
@@ 2.2 PDHG
PDHG serves as the base routine of cuPDLPx to solve the primal-dual problem (2). Specifically, the update rule of PDHG on (2) is given as:
@@
where $\sigma$ and $\tau$ are dual and primal stepsizes respectively. cuPDLPx further reparameterizes as $\tau=\frac{\eta}{\omega}$ and $\sigma=\eta\omega$, where $\eta$ is called the stepsize and $\omega$ is the primal weight. And the canonical norm of PDHG is defined by $\|\cdot\|_{P}$, where $P:=P_{\eta,\omega}=\begin{bmatrix}\frac{\omega}{\eta}I&A^{\top}\\ A&\frac{1}{\eta\omega}I\end{bmatrix}$. For notational convenience, we define the primal-dual iterate at iteration $k$ as $z^{k}=\begin{bmatrix}x^{k}\\ y^{k}\end{bmatrix}$, and write $z^{k+1}=\operatorname{PDHG}(z^{k})$ to denote a single PDHG update step applied to $z^{k}$ as defined in (3).
@@ 3 Algorithmic enhancements on top of cuPDLP
In this section, we discuss a series of algorithmic further enhancements used in cuPDLPx to improve the efficiency and convergence behavior of the algorithm, on top of what is used in cuPDLP . These enhancements are primarily motivated by recent theoretical advances in restarted Halpern PDHG, as described in [16].
At the high-level, cuPDLPx is based on restarted Halpern PDHG algorithm (in contrast to the restarted average PDHG algorithm used in cuPDLP ), and includes additional techniques such as the reflected updates, adaptive restarting, stepsize selection and dynamic primal-dual weight. Each component is carefully designed and implemented to accelerate convergence and enhance stability for solving linear programming problems. We describe these contributions below.
Halpern scheme. The Halpern scheme, along with its reflected variant, is an enhancement to the PDHG algorithm that improves convergence properties and accelerates solver performance. Originally developed to accelerate general operator splitting methods, the Halpern scheme has been successfully adapted to linear programming with strong theoretical guarantees and improved preliminary numerical performance [16]. Halpern PDHG interpolates between the current PDHG iterate and the initial point, using a weighted average. Specifically, the update rule at iteration $k$ is:
@@
As discussed in [16], the Halpern and average schemes share structural similarities but differ in theoretical guarantees. Notably, the Halpern scheme supports reflection, which enables the use of larger stepsizes and often leads to faster empirical convergence, as described below.
Reflection. The reflection technique builds upon the Halpern scheme by applying a reflected version of the PDHG operator, $(1+\gamma)\operatorname{PDHG}-\gamma\operatorname{id}$ with reflection $\gamma\in[0,1]$, instead of the vanilla PDHG update. This is also called over-relaxation in the literature of variational inequality. The resulting update rule of the Reflected Halpern PDHG update is
@@
This reflection mechanism effectively allows for longer steps than the vanilla Halpern update and has been shown, both theoretically and empirically, to improve convergence [16]. It is one of the key drivers of cuPDLPx ’s superior performance compared to cuPDLP .
Adaptive restart. Restarting is a key enhancement on first-order methods for attaining high-accuracy solutions for linear programming [6]. A restart is triggered when certain progress criteria are met, and the algorithm restarts from a new initial solution. In the Halpern scheme, the anchor is periodically reset to the current solution. This helps the algorithm stay focused on the neighborhood of the optimal solution, especially as the initial anchor becomes increasingly outdated during iterations. Such restart strategies have been shown to accelerate convergence rates for Halpern PDHG [16] in theory.
cuPDLPx adopts an adaptive restart strategy that evaluates potential restart conditions at each iteration. The core idea is to monitor a fixed-point error metric $r(z)=\|z-\operatorname{PDHG}(z)\|_{P}$ at solution $z$, and trigger a restart when specific decay patterns are observed. This fixed-point error metric is motivated by the recent theoretical insight [16], which shows that this fixed-point error is a natural metric for the restarted Halpern PDHG. It is different from the normalized duality gap used in CPU-based PDLP [4, 5] and the KKT error used in cuPDLP [15].
Three restart conditions are used:
@@
This combined restart strategy balances aggressiveness and stability, contributing to cuPDLPx’s robust and accelerated convergence across a wide range of linear programming instances.
Stepsize. In contrast to an adaptive stepsize heuristic used in cuPDLP, cuPDLPx switches to constant stepsize $\eta=\frac{0.998}{\|A\|_{2}}$, where operator norm of constraint matrix $\|A\|_{2}$ is approximated by power iteration. The rationale behind this change is twofold. First, the reflected Halpern PDHG inherently selects a larger effective stepsize, and the adaptive stepsize strategy from [15] can become overly aggressive in certain instances, leading to occasional instability. Second, a constant stepsize eliminates the need for the sequential stepsize search, which is not perfectly-suited for parallel execution on GPUs. Thus, this modification improves both stability and parallel efficiency in cuPDLPx.
Primal weight update. The primal weight $\omega$ is a crucial parameter in balancing progress between the primal and dual spaces. Our empirical experiences have shown that convergence behavior in PDLP is highly sensitive to the choice of this weight. The intuition for adjusting the primal weight $\omega^{n}$ is to set it such that the weighted distances to optimality in the primal and dual domains are in the same scale, i.e., $\|(x^{n,t}-x^{*}),0\|_{w^{n}}\approx\|(0,y^{n,t}-y^{*})\|_{w^{n}}$. However, the optimal solution $(x^{*},y^{*})$ is not known during the iterations.
To address this, cuPDLP estimates the primal and dual distances to optimality based on the observed movements in the previous epoch. The distance estimates are smoothed using exponential averaging and subsequently used to update the primal weight at each restart.
In cuPDLPx , we further enhance this strategy by modeling the update as a control problem and introducing a PID controller to dynamically regulate the primal weight. We define the error as the gap between the primal and dual distances on a logarithmic scale:
$e^{n}=\log\left(\frac{\sqrt{w^{n}}\left\|x^{n,t}-x^{n,0}\right\|_{2}}{\frac{1}{\sqrt{w^{n}}}\left\|y^{n,t}-y^{n,0}\right\|_{2}}\right)$
@@
This PID-controlled update ensures that the primal-dual balance is continuously adjusted in a stable manner, improving convergence speed and robustness across a wide range of problem instances.